solve the proportion 5/7 = 20/d
Answer by haileytucki(390)

(5)/(7)=(20)/(d)
Since d is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
(20)/(d)=(5)/(7)
Since there is one rational expression on each side of the equation, this can be solved as a ratio. For example, (A)/(B)=(C)/(D) is equivalent to A*D=B*C.
20*7=5*d
Since d is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
5*d=20*7
Multiply 5 by d to get 5d.
5d=20*7
Multiply 20 by 7 to get 140.
5d=140
Divide each term in the equation by 5.
(5d)/(5)=(140)/(5)
Simplify the left-hand side of the equation by canceling the common factors.
d=(140)/(5)
Simplify the right-hand side of the equation by simplifying each term.
d=28